OpenStudy (anonymous):

Choose the correct simplification of 10 over x to the power of negative 9. 10x9 x to the 9th power over 10 10 over x to the 9th power

OpenStudy (anonymous):

10x^9

OpenStudy (anonymous):

should be 10 over x to the 9power i believe, that is if the original problem is (10)/(x^-9)

OpenStudy (anonymous):

is it like ths : \[10/(x)^-9

OpenStudy (anonymous):

yh i was thinking maybe it was (10/x)-9

OpenStudy (anonymous):

oh sorry : \\[10/(x)^-9\]

OpenStudy (anonymous):

but then no answer for that given so

OpenStudy (anonymous):

it is c

OpenStudy (anonymous):

now see here x^-a = 1/x^a now here it would be 10 * x^ 9 = 10(x)^9

OpenStudy (anonymous):

if the problem is (10)/(x^-9), the only possible simplification is : (10)/(x^9)